OpenStudy (anonymous):

Use the substitution method to solve this system of equations: y = 3x + 8 2x + 6y = 14

hartnn (hartnn):

so substitute y= 3x+8 in your 2nd equation \(\large 2x +6(3x+8) =14\) can you solve this for x ?

OpenStudy (anonymous):

2x + 18x + 48 = 14 20x + 48 = 14 -48 = -48 20x = -34 20x/20 = -34/20 x = -1.7

hartnn (hartnn):

thats correct! :) now find 'y' y =3x+8 = 3(-1.7) + 8 =...

OpenStudy (anonymous):

y = -5.1 + 8 y = 2.9

hartnn (hartnn):

thats correct too! good work :)

OpenStudy (anonymous):

Thank you so much :D

hartnn (hartnn):

most welcome ^_^